The hypotenuse of a right triangle is 70 ft and one leg is 25 ft. Find the length of the other side.
Answer by checkley77(12844)

X^2+25^2=70^2
X^2+625=4900
X^2=4900-625
X^2=4275
X=SQRT4275
X=65.38 ANSWER.